### v2

# Hybrid Machine-Learning Particle Identification for the ePIC Proximity-Focusing RICH

Abstract
We present a machine-learning–based particle-identification study for the proximity-focusing Ring Imaging Cherenkov (pfRICH) detector of the ePIC experiment at the Electron–Ion Collider. Operating in the backward region ($3.5\lesssim\eta\lesssim-1.5$), the pfRICH is designed to provide at least $3\sigma$ separation among pions, kaons, and protons up to 7 GeV/$c$ for Semi-Inclusive Deep Inelastic Scattering measurements. Using a standalone Geant4 simulation of the pfRICH, we develop a hybrid model that combines convolutional neural network–based feature extraction with gradient-boosted decision-tree classifiers. This approach significantly improves Cherenkov-ring pattern recognition and particle separation performance, demonstrating the potential of hybrid machine-learning techniques for next-generation Cherenkov detectors at the EIC.
1 Introduction
The proximity-focusing Ring Imaging Cherenkov (pfRICH) serves as a Particle Identification (PID) detector in the backward region, $3.5\lesssim\eta\lesssim-1.5$, of the upcoming ePIC experiment at the Electron–Ion Collider (EIC). The ability to distinguish final-state hadron production in electron–nucleus collisions is crucial for Semi-Inclusive Deep-Inelastic Scattering (SIDIS) measurements. The pfRICH consists of three main components: an aerogel radiator, a mirror system, and a photon sensor plane as illustrated in Fig. 1.
Figure 1: Schematic view of the proximity-focusing Ring Imaging Cherenkov (pfRICH) detector.
Charged particles traversing the aerogel emit Cherenkov photons, which are reflected by mirrors and projected onto the photon sensor plane, forming characteristic ring patterns. The measurement of the ring radius allows reconstruction of the Cherenkov angle, which is directly related to the velocity of the particle. The Cherenkov emission angle is defined as
$\cos{\left(\theta_{c}\right)}=\frac{1}{n\beta},$ (1.1)
where $n$ is the refractive index of the aerogel and $\beta=\frac{v}{c}$ is the particle’s velocity in units of the speed of light. For relativistic particles ($\beta\lesssim 1$), expanding to first order in $m^{2}/p^{2}$ gives
$\theta_{c}^{\,2}\simeq\theta_{\mathrm{sat}}^{\,2}-\frac{1}{n}\frac{m^{2}}{p^{2}},$ (1.2)
where $\theta_{\mathrm{sat}}$ is the saturation angle corresponding to ultra-relativistic emission.
This implies that for a given momentum measured by the tracking system, the Cherenkov angle $\theta_{c}$ depends on the particle mass (see Fig. 2). This property allows the pfRICH detector to separate pions, kaons, and protons by comparing the measured angle with the expected values. The pfRICH is expected to provide at least $3\sigma$ separation up to 7 GeV/$c$ [Chatterjee:2024zrn].
Figure 2: Left: Reconstructed Cherenkov angle (mrad) as a function of hadron’s momentum (GeV/c) for electrons, pions, kaons and protons using Eq. (1.1). Right: Overlapping reconstructed rings of pion and kaon particles on the sensor plane.
Beyond standard reconstruction methods, machine-learning techniques are explored in order to improve the pfRICH particle-identification performance. The Cherenkov ring patterns recorded on the photon sensor plane form a natural image-like structure that can be exploited by using advanced pattern-recognition algorithms. In the next section, we introduce a hybrid model that combines pattern recognition with boosted decision trees, using both hadron kinematic information and hit-level features to improve pfRICH PID performance.
2 Simulation and data
The datasets used in this study are produced using a standalone Geant4-based pfRICH simulation [eic_pfrich_v110], which provides a detailed description of the aerogel radiator, mirror system, and photosensor response. Single-particle samples are generated for electrons ($e^{-}$), pions ($\pi^{-}$), kaons ($K^{-}$) and protons ($p$) over a momentum range of 111212 GeV/$c$ and pseudorapidities $3.5\leq\eta\leq-1.5$, with uniformly distributed azimuthal angles. The simulation is performed on a structured $(p,\eta)$ grid to ensure uniform coverage of the pfRICH acceptance.
To prepare the raw detector output for machine-learning training, a sanitization and quality-assurance (QA) procedure is applied to the Geant4 hit records. Only primary particles corresponding to the target species ($e^{-}$, $\pi^{-}$, $K^{-}$, $p$) are retained in order to suppress secondary particles originating from material interactions. Hits with energies below 1 eV and those outside the active fiducial sensor area are removed to reject optical background and dark noise. The resulting sanitized hitmaps (Fig. 3) isolate the primary Cherenkov signal and provide clean, sparse patterns suitable for pattern-recognition models.
Figure 3: Comparison of photon hit patterns on the pfRICH sensor plane before and after preprocessing. Left: raw event, including background contributions. Right: corresponding sanitized event used for machine-learning training. The sanitization procedure removes low-energy hits (<1<1 eV) and applies fiducial cuts to isolate the primary Cherenkov signal.
A radial acceptance cut of $R<650$ mm is applied to preselect hits within the active sensor plane. The final sanitized datasets contain 2,260,869 events distributed across the four particle species: 625,812 electrons, 568,140 pions, 528,816 kaons, and 538,101 protons.
Quality-assurance distributions of the sanitized datasets are presented in Fig. 4. The $x$ and $y$ coordinate distributions (top panels) exhibit the segmented structure of the photosensor array, with gaps corresponding to inactive regions between sensor tiles. The radial distribution (bottom left) shows distinct peaks associated with Cherenkov rings at different momenta, with all hits contained within the $R_{\mathrm{max}}=650$ mm acceptance. The two-dimensional hitmap (bottom right) displays the integrated ring pattern on the pfRICH sensor plane, including the central aperture corresponding to the beam pipe.
Figure 4: Quality-assurance distributions for sanitized photon hits. Top: $x$ and $y$ coordinate distributions showing the segmented photosensor structure. Bottom left: radial distribution with peaks corresponding to Cherenkov rings at various momenta; the red dashed line indicates the $R_{\mathrm{max}}=650$ $m$ acceptance cut. Bottom right: two-dimensional hitmap on the pfRICH sensor plane showing the integrated ring pattern.
The aggregated hit patterns separated by particle species are shown in Fig. 5. Electrons ($e^{-}$) produce the largest rings, corresponding to the saturation angle $\theta_{\mathrm{sat}}$ since $\beta\approx 1$ across the full momentum range. Pions, kaons, and protons produce progressively smaller rings at the same momentum, reflecting their larger masses and correspondingly smaller Cherenkov angles according to Eq. (1.2). The clear ring structure observed in these aggregated hitmaps motivates the use of geometric features, such as the ring radius, together with pattern-recognition techniques for particle identification.
Figure 5: Aggregated hit maps on the pfRICH sensor plane for each particle species. Top left: electrons. Top right: pions. Bottom left: kaons. Bottom right: protons. The characteristic Cherenkov ring radius decreases with increasing particle mass at fixed momentum, providing the physical basis for particle identification. The red dashed circle indicates the $R_{\mathrm{max}}=650$ $m$ acceptance boundary.
The datasets are divided into training and test sets using a stratified random split at the event level. This stratification guarantees that each particle species ($e^{-}$, $\pi^{-}$, $K^{-}$, $p$) is proportionally represented in both subsets, thereby preventing class imbalance from biasing the evaluation. A fixed random seed is used to ensure reproducibility.
3 Machine learning model
The hybrid model combines a convolutional neural network (CNN) encoder with an XGBoost gradient-boosted decision-tree classifier [Chen_2016]. The pipeline consists of two main stages:
  1. 1.
    a CNN encoder for extracting spatial features from per-event photon hit patterns, and
  2. 2.
    an XGBoost classifier for particle identification.
Together, these components form a multi-stage pipeline optimized for combining image-like detector information with tabular kinematic inputs, as illustrated in Fig. 6.
I. Per-event histogram construction Build $64\times 64$ TH2 images from photon hitsII. CNN feature extraction Encode images $\rightarrow$ latent embeddings $\mathbf{z}$III. Feature combination Concatenate $\mathbf{z}$ with beam particle featuresIV. XGBoost training $f_{\mathrm{xgb}}(\mathbf{z},\mathbf{x}_{\mathrm{beam}})$V. Inference pipeline Histogram $\rightarrow$ CNN $\rightarrow$ XGBoost $\rightarrow$ PID
Figure 6: Algorithm flow of the hybrid CNN+XGBoost model, illustrating per-event histogram construction, CNN feature extraction, feature combination, and classifier inference.
3.1 Pattern recognition
The CNN encoder processes per-event photon hit patterns to extract spatial features that capture the geometric structure of Cherenkov rings. For each event, photon hits are binned into a $64\times 64$ histogram representing the spatial distribution on the sensor plane. The CNN architecture consists of:
  1. 1.
    Convolutional layers: three Conv2D layers with 32, 64, and 128 filters, using ReLU activation and max-pooling to extract hierarchical spatial features;
  2. 2.
    Global average pooling: reducing the spatial dimensionality while preserving the learned feature content;
  3. 3.
    Dense projection: mapping the pooled features to a 64-dimensional latent embedding $\mathbf{z}$ with layer normalization.
The CNN encoder is not trained end-to-end with the classifier. Instead, the latent embeddings are computed separately and subsequently combined with kinematic features for classification.
3.2 XGBoost
The XGBoost classifier operates on a 72-dimensional input vector, composed of eight kinematic and detector-level features augmented by the 64-dimensional CNN latent representation. The kinematic and detector features include the transverse momentum $p_{T}$, total momentum magnitude $P_{\mathrm{total}}=|\vec{p}|$, pseudorapidity $\eta$, polar and azimuthal angles $(\theta_{\mathrm{beam}},\phi)$, the reconstructed Cherenkov angle $\theta_{c}$, the mean photon arrival time $T_{\mathrm{hits}}$, and the number of detected photon hits $n_{\mathrm{hits}}$. These quantities are concatenated with the CNN embedding $\mathbf{z}$ to form the input to the gradient-boosted classifier.
4 Results
The performance of the hybrid CNN+XGBoost model is evaluated on an independent test set of 570,963 events, while 1,712,886 events are used for training. The classifier operates on a 72-dimensional feature vector obtained by combining 64-dimensional CNN embeddings extracted from per-event hit patterns with eight beam particle features ($P_{\mathrm{total}}$, $p_{T}$, $\eta$, $\theta_{\mathrm{beam}}$, $\phi$, reconstructed Cherenkov angle, mean photon arrival time, and number of detected photon hits).
4.1 Model performance
The XGBoost classifier is trained using 300 boosting rounds, a maximum tree depth of 6, and a learning rate of 0.1. Table 1 summarizes the overall classification performance of the hybrid model.
Model Accuracy F1-score Precision Recall
CNN+XGBoost 95.07% 0.951 0.951 0.951
Table 1: Overall classification metrics for the hybrid CNN+XGBoost model using combined CNN embeddings and beam particle features.
Table 2 reports the per-class identification performance. Protons achieve the highest efficiency (98.4%), reflecting their distinct Cherenkov signature at lower $\beta$ values. Kaons, which are particularly challenging for traditional reconstruction due to their intermediate mass, reach an efficiency of 97.2% with the hybrid approach. Electrons and pions achieve efficiencies of 93.7% and 94.6%, respectively.
Particle Precision Recall F1-score
$e^{-}$ 0.94 0.94 0.94
$\pi^{-}$ 0.95 0.95 0.95
$K^{-}$ 0.97 0.97 0.97
pp 0.98 0.98 0.98
Table 2: Per-class classification metrics for the hybrid CNN+XGBoost model.
4.2 Momentum dependence of particle identification performance
The momentum dependence of the classifier performance is essential for assessing the physics validity of the model. Figure 7 shows the identification efficiency as a function of the true particle momentum for each species. The integrated efficiencies are 93.7% for electrons, 94.6% for pions, 97.2% for kaons, and 98.4% for protons.
The efficiency of identifying a given particle species $i$ is defined as
$\varepsilon_{i}=\frac{N(\hat{y}=i,\,y=i)}{N(y=i)}=P(\hat{y}=i\mid y=i),\qquad i\in\{e,\ \pi,\ K,\ p\},$ (4.1)
where $y$ denotes the true particle species and $\hat{y}$ the species predicted by the classifier. The efficiency therefore quantifies the probability of correctly identifying a given particle $i$ in the presence of the competing hypotheses $\{e,\pi,K,p\}$.
Figure 7: PID efficiency versus momentum for the CNN+XGBoost model. Efficiencies remain above 95% for kaons and protons over the full momentum range, while electrons and pions exhibit a moderate degradation at high momentum.
The mis-identification probabilities are presented in Fig. 8 as a function of momentum for pion, kaon, and electron truth samples. For electrons, the dominant confusion channel is $e\to\pi$, which increases with momentum as Cherenkov angles converge in the ultra-relativistic regime, while $e\to K$ remains suppressed at low momentum and rises only at high momentum. Kaon mis-identification is primarily driven by confusion with pions at intermediate and high momentum, whereas the $K\to p$ channel remains subdominant. For pions, the leading confusion channel is $\pi\to K$, which increases with momentum, while $\pi\to e$ remains at the percent level and $\pi\to p$ is strongly suppressed. Overall, mis-identification probabilities remain well below the 10% level up to 12 GeV/$c$, indicating robust PID performance in the high-momentum regime.
Figure 8: Mis-identification probabilities as a function of momentum for (a) pion, (b) kaon, and (c) electron truth samples. Only the dominant confusion channels are shown.
4.3 Separation power
The separation power between two particle species $(i,j)$ is quantified in terms of an effective number of standard deviations, $n_{\sigma}$. For a given momentum bin, $n_{\sigma}$ is defined as [Ghosh_2025]
$n_{\sigma}=\frac{|\mu_{i}-\mu_{j}|}{\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}},$ (4.2)
where $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the classifier score for species $i$. Dedicated binary XGBoost classifiers are trained separately for $\pi/K$ and $e/\pi$ separation to evaluate the momentum dependence of the separation power.
The $e/\pi$ and $\pi/K$ separation power as a function of momentum is shown in Fig. 9. Here, the hybrid model achieves separation exceeding $3\sigma$ up to approximately 8 GeV/$c$. At higher momenta, the separation gradually decreases as particle velocities converge, while remaining slightly above $3\sigma$ up to about 12 GeV/$c$.
Figure 9: Separation power as a function of momentum for $\pi/K$ and $\pi/e$, evaluated using dedicated binary classifiers. In both cases, the separation decreases at high momentum as particle velocities converge ($\beta\to 1$), while remaining above $3\sigma$ (dotted line) up to 12 GeV/$c$.
4.4 Feature importance analysis
The feature-importance in Fig. 10 shows the XGBoost classifier rankings. The reconstructed Cherenkov angle $\theta_{c}$ dominates the ranking, consistent with its direct physical relationship to particle mass through Eq. (1.2). The total momentum $P_{\mathrm{total}}$ ranks second, followed by the pseudorapidity $\eta$ and transverse momentum $p_{T}$.
Figure 10: Feature-importance ranking for the hybrid CNN+XGBoost model. Features are color-coded by type: beam/kinematic features (blue) and CNN latent features (orange).
Notably, several CNN latent features appear among the top 20 most important features, demonstrating that the CNN encoder extracts complementary information from hit patterns that is not fully captured by the tabular beam features alone, thereby validating the hybrid approach. To further assess interpretability, the CNN latent space is diagonalized using a principal-component basis, and the feature-importance hierarchy is found to be stable, with beam kinematics and the reconstructed Cherenkov angle remaining dominant, while only the leading CNN mode contributes at a sub-dominant level.
5 Conclusion
We have presented a hybrid machine-learning approach to enhance particle-identification (PID) performance for the proximity-focusing RICH detector of the ePIC experiment at the Electron–Ion Collider. Efficient hadron separation over a broad momentum range in the backward region is a key requirement for precision measurements of hadronic structure in electron–nucleus collisions.
Using a realistic standalone Geant4 simulation of the pfRICH, we have demonstrated that machine-learning techniques can significantly improve the reconstruction of Cherenkov patterns and the discrimination between pions, kaons, and protons across the relevant kinematic phase space. The proposed hybrid model, combining CNN-based pattern recognition with gradient-boosted decision trees, achieves separation power exceeding $3\sigma$ up to 12 GeV/$c$ for both $e/\pi$ and $\pi/K$ using dedicated binary classifiers, and satisfies the pfRICH design requirements.
These results indicate that advanced ML-based PID methods can reinforce and extend the physics capabilities of the EIC [AbdulKhalek:2021gbh]. The approach is readily applicable to other ePIC sub-detector systems and provides a flexible framework for integrating image-based detector information with traditional kinematic observables.
As a next step, the model will be evaluated within the full ePIC software framework, including realistic background conditions. Future developments will also explore dedicated denoising techniques to further enhance signal quality and strengthen the PID performance of the pfRICH and related detector systems.
Acknowledgments
We thank the ePIC collaboration for providing the pfRICH detector simulation framework and for valuable discussions related to particle-identification performance at the Electron–Ion Collider. We are grateful to the developers of the standalone Geant4 pfRICH simulation for making their tools publicly available.
This work is supported by the Simons Foundation.